# ICP-4D: Bridging Iterative Closest Point and LiDAR Panoptic Segmentation

### v1

@@ 3.2.2 Dynamic Instance
Returning to the ICP procedure, we follow the standard process, initializing transformations with a histogram-based method [26] to reduce sensitivity to initial estimates and subsequently running iterative refinement to get the final rigid transformation. After obtaining the transformation, we determine whether two instances correspond by evaluating the IoU between their transformed point sets. Specifically, points with errors smaller than the predefined threshold $\tau_{dist}$ are treated as inliers. Then, in order to filter valid matching candidates, we evaluate alignment accuracy between source and destination instances by IoU scores as follows:
$\mathcal{M}^{dyn}=\begin{cases}\text{matched},&\!\!\!\!\text{if}~\langle\mathcal{F}(P(\mathcal{G}^{\prime\prime}_{\text{src}})),P(\mathcal{G}^{\prime\prime}_{\text{dst}})\rangle\geq\tau_{\text{iou}},\\ \text{rejected},&\!\!\!\!\text{otherwise},\end{cases}$ (8)
where $\langle\cdot\rangle$ denotes the IoU calculation, and $\mathcal{F}$ applies the estimated transformation obtained by ICP.
Option: Bijective assignment. A naive approach is to pair each source instance with its most similar target instance; however, it fails to enforce one-to-one correspondence. Thus, we derive a transformation-aware matching cost that integrates the ICP results—rotation $\mathbf{R}$, translation $\mathrm{t}$, and the IoU score $s_{\text{iou}}$. Formally, the matching cost $\mathcal{C}$ of dynamic matching pairs $(i,j)\in\mathcal{M}^{dyn}$ is defined as:
$c_{ij}^{(t)}=\|\mathrm{t}_{ij}\|_{2},$ $c_{ij}^{(r)}=\|\theta_{ij}\|_{2},\quad c_{ij}^{(s)}=\|1-s_{iou}\|_{2},$
where $\theta$ denotes the axis–angle rotation vector obtained from $\mathbf{R}$. We normalize each cost term to comparable scales before weighting so that $\gamma_{t}$, $\gamma_{r}$, and $\gamma_{s}$ control their relative contributions. By minimizing $\mathcal{C}$ with the Hungarian algorithm [21], we obtain a bijective assignment that considers the geometric alignment.
3.2.3 Missing Instance
Occlusions stemming from the inherent limitations of LiDAR sensor often break temporal continuity beyond two consecutive scans. To mitigate this, we introduce a memory bank $\mathcal{B}$ that caches the world coordinates, semantic labels, and instance IDs of unmatched objects. Owing to bounded memory capacity, $\mathcal{B}$ keeps only entries observed within a short window of the previous $w_{\text{mem}}$ scans. If $\mathcal{B}$ is nonempty, we retrieve previously stored instance information and associate using the dynamic instance matching process.
@@ 4 Experiments
Table 1: Quantitative results on SemanticKITTI benchmark [4]. We highlight the best scores in bold and the underline indicates the second-best scores. Note that * represents that the results are obtained from our implementation following the official training setup.
@@
4D-PLS [2] 4 62.7 65.1 60.5 65.4 61.3 56.9 56.4 57.4 66.9 51.6
4D-StOP [20] 4 67.0 74.4 60.3 65.3 60.9 63.9 69.5 58.8 67.7 53.8
Eq-4D-StOP [63] 4 70.1 77.6 63.4 66.4 67.1 67.0 72.0 62.4 69.1 60.9
Mask4Former [54] 2 70.5 74.3 66.9 67.1 66.6 68.4 67.3 69.6 72.7 65.3
@@
KPConv [43]+PP [23]+MOT [47] 1 46.3 37.6 57.0 64.2 54.1 38.0 25.9 55.9 66.9 47.7
KPConv+PP+SFP [33] 1 46.0 37.1 57.0 64.2 54.1 38.5 26.6 55.9 66.9 47.7
black!10black DS-Net[15] 1 - - - 63.5 64.6 62.0 - - 60.6 66.9 52.0
@@
Table 2: Quantitative results on panoptic nuScenes [13].
@@
4D-StOP [20] 4 60.5 62.5 58.6 74.4 54.9
Eq-4D-StOP [63] 4 67.3 73.7 61.5 76.4 58.7
Mask4Former [54] 2 67.6 63.0 72.5 76.6 66.3
black!10black       (b) Query-propagated Association
Mask4D [30] 1 58.5 54.1 63.1 75.5 55.7
black!10black       (c) Detect & Track Association
@@
[…]
4.1 Experiment Setup
Datasets. We choose two representative benchmarks. SemanticKITTI [4] captures point clouds using a 64-beam LiDAR at 10 Hz and provides point-wise annotations for 19 semantic classes (8 Things / 11 Stuff). In contrast, panoptic nuScenes [13] uses a 32-beam LiDAR operating at 2 Hz and offers point-wise annotations for 16 semantic classes (10 Things / 6 Stuff). This contrast serves as a valuable testbed for assessing generalization across dense (SemanticKITTI) and sparse (nuScenes) LiDAR settings.
Evaluation Metrics. To evaluate both tracking and segmentation quality, we adopt the widely used LSTQ metric [2], defined as $\text{LSTQ}=\sqrt{S_{cls}\times S_{assoc}}$. Here, $S_{cls}$ measures how accurately semantic labels are assigned and $S_{assoc}$ evaluates how consistently instance identities are maintained throughout the sequence, independently of semantic accuracy. However, $S_{assoc}$ is computed only from true positive associations weighted by IoU scores. Thus, when prediction masks are stable and yield high IoU scores, ID switch causes a sharper drop in this score. Consequently, while ICP-4D primarily influences $S_{assoc}$, the LSTQ must be taken into account as a meaningful indicator of temporal consistency.
Implementation Details. Our proposed ICP-4D can be incorporated into any 3D LiDAR panoptic segmentation network. For fair comparison, we basically follow two baseline’s configuration, DS-Net [15] and Mask4Former [54]. Since Mask4Former is originally developed for 4D panoptic segmentation, we train it using the official settings with the input modified to a single scan instead of multiple sweeps. For instance association, thresholds for filtering instance pairs are set to $\tau_{center}=\tau_{cov}=\tau_{dist}=0.1$ and $\tau_{iou}=0.2$. Additionally, the short window $w_{\text{mem}}$ is set to $3$ for missing instances, the regularization parameter for optimal transport is set to $\epsilon=0.2$, and the number of ICP iterations is fixed to 30. All experiments are conducted on a single RTX A6000.
Figure 3: Qualitative comparison on SemanticKITTI validation set. We visualize the association results over five consecutive scans for both the baselines and our ICP-4D. Different colors represent different instances. The dotted boxes zoom in for clear comparison. For clarity, we use circled numbers (①–⑤) to represent the frame indices.
Figure 4: Comparison of computational efficiency. We illustrate the efficiency trade-offs between memory usage (left) & runtime (right) and performance across different methods. Yellow triangles show LSTQ scores for each model. Red dashed line denotes our memory & runtime, and yellow dashed line marks our LSTQ score.
4.2 Quantitative Analysis
4.2.1 4D LiDAR Panoptic Segmentation Results
SemanticKITTI. Table 4 reports the experimental results on SemanticKITTI validation and hidden test set. Here, we clearly observe two insights. First, ICP-4D achieves the highest LSTQ score across all data split regardless of the number of input scans. Specifically, while (a) IoU-based association approaches rely on 22-44 more input point clouds, our ICP-4D achieves superior performance using only a single scan, showing improvements of up to 13.2 percent points in $S_{assoc}$ and 10.0 percent points in LSTQ. Furthermore, by comparing with the learning-based approaches (b) and (c), another noticeable aspect is that ICP-4D consistently outperforms recent state-of-the-art methods on both the validation and test sets, whereas Mask4D and CA-Net drop in performance on the test set. This highlights the robustness of our direct geometric registration with raw point clouds, rather than relying on training-dependent feature association.
@@ 4.2 Quantitative Analysis
nuScenes. Table 4 shows the 4D panoptic results on panoptic nuScenes dataset. Consistent with the observations on the SemanticKITTI dataset, ICP-4D delivers notably stronger instance association performance, outperforming existing methods in terms of LSTQ. It is important to note that baselines generally suffer when the LiDAR scans are captured with a long interval. In such scenarios, multi-scan aggregation benefits from denser temporal support and thus exhibit strong performance. For example, EQ-4D-StOP with 4 aggregated scans surpasses Mask4Former with 2 scans by a substantial margin in the association metric. Nevertheless, our ICP-4D demonstrates that even without accessing multiple scans, a single-scan input is sufficient to achieve competitive performance. This result implies that ICP-4D remains highly effective even under sparse LiDAR configurations.
4.2.2 Comparison of Efficiency
@@
Next, we examine computational efficiency in terms of memory usage and runtime. For each method, we report the peak GPU memory consumption and the average runtime measured over the entire validation set. Figure 4 illustrates the results, supporting that ICP-4D offers a most well-balanced trade-off between performance and computational overhead. First, when compared with the IoU-based association approaches, we observe that ICP-4D requires 60.7%–79.4% less memory and achieves a 26.1%–45.6% runtime efficiency improvement, all while delivering superior LSTQ performance. While Mask4D exhibits high efficiency in runtime, it inevitably incurs substantial memory overhead due to its transformer-based architecture design. Lastly, CA-Net is computationally efficient; however, it fails to obtain reliable association performance. This gap is already observable on the validation set and becomes even more pronounced on the hidden test set, as highlighted in blue in Table 4.
4.3 Qualitative Analysis
@@
In this section, we deeply analyze the effectiveness of ICP-4D qualitatively by presenting the visualization results in Figure 3. In the first row, there are two bicyclists riding in a line. Both ground truth and our ICP-4D consistently distinguish the two bicyclists across time. However, Mask4Former struggles to preserve identity consistency when objects are spatially adjacent or overlapping due to its IoU-based association. Furthermore, Mask4D employs learnable tracking queries for association, which can lead to confused tracking queries when objects share similar visual or spatial features. Then, in the second row, we emphasize the effectiveness of our memory bank in handling missing instances. Even when the 3D panoptic model fails to detect the person colored in green at the second time step, the memory bank successfully retrieves the missing instance and preserves its identity across scans. In contrast, Mask4Former experiences ID switching when it fails to detect instance due to the absence of the overlapped region. Consequently, these observations underscore the importance of explicit geometric matching in our approach and a well-structured association pipeline.
@@
4.4 Ablation Study
Module Design. Table 3 shows the effects of our proposed methods. Exp. #1 directly applies the vanilla Iterative Closest Point (ICP) for associating instances, serving as a baseline to examine the feasibility of ICP in an unsupervised manner. Despite its strong matching capability, ICP is highly sensitive to the quality of initial alignment. In Exp. #2, we observe a little performance gain when applying histogram-based initialization proposed by [26], which stabilizes the initial alignment. Notably, our Sinkhorn-based soft matching brings further improvements without the influence of initial pose (see Exp. #3). The key takeaway is that instance-aware correspondence to alleviate the instability of local point matching is highly beneficial for obtaining reliable matching points between noisy instance segments, and as a result helps to estimate an accurate transformation. Exp. #4 demonstrates that coupling these components yields a complementary effect, leading to more robust instance association and improved overall performance. Lastly, Exp. #5 demonstrates the effectiveness of incorporating a memory bank for handling occluded and missing instances, resulting in superior performance.
Effect of Different Types of Instances. To validate our carefully designed pipeline with respect to different instance types, Table 4 presents the association results and average inference time per a single scan. First, we employ the association strategy designed for static instances (Sec. 3.2.1) to matching candidates over all detected objects. Despite requiring minimal computation, only a small subset of instances remain consistently tracked, 18.5 percent points indicating that static cues alone are insufficient for reliable association. In contrast, relying solely on dynamic instance association (Sec. 3.2.2) achieves high performance, but is computationally expensive. Thus, combining these strategies reduces the runtime by 48.2% while simultaneously improving performance, as it significantly narrows the matching candidates for dynamic instance association. Consequently, our full ICP-4D enhances the tracking continuity without sacrificing computational efficiency.
Table 5: Ablation on matching strategy.
Method LSTQ $_{\text{assoc}} S$$_{\text{cls}}$
Hungarian Matching 72.3 77.6 67.5
Greedy Matching 72.7 78.3 67.5
Table 6: Ablation on DBSCAN instance refinement.
@@ 4.4 Ablation Study
Analysis of DBSCAN Instance Refinement. In this specific experiment of integrating ICP-4D with Mask4Former, we apply DBSCAN clustering [10] as done in the original Mask4Former setting. Table 6 summarizes the results, where we derive two major insights. First, ICP-4D consistently surpasses existing state-of-the-art methods, both with and without the additional clustering step, achieving at least a 0.4 percentage point gain in LSTQ. Furthermore, we discover that the increase of runtime primarily results from the post-processing step. Without this step, ICP-4D delivers the most efficient runtime of 840 ms compared to all other methods, with only a marginal loss in performance. These observations clearly highlight that ICP-4D offers both high efficiency and strong performance, and can further benefit from improvements in the backbone network.
[…]
Effect of Sinkhorn-based Soft Matching. The Sinkhorn-based soft matching forms instance-aware point correspondence between source and target points. Owing to this strategy, it allows accurate error measurements albeit with outliers, by not being overwhelmed by the local proximity. Figure 5 visualizes the point-wise correspondences obtained from two consecutive scans, with the current scan serving as the source and the previous scan as the target. For this analysis, we extract each point segment at the initial stage of ICP, as this stage retains its geometric structure intact. As shown in the highlighted region, the nearest neighbor matching is biased toward the closest point, ignoring the underlying point-wise relationships of the instance. In contrast, applying Sinkhorn-based soft matching on the baseline maintains the geometric structure, yielding correspondences that align consistently with the instance’s motion direction.
Class-wise Performance Analysis. In this section, we further analyze the effectiveness of ICP-4D w.r.t. the class over the SemanticKITTI validation set. We categorize classes with a large number of observed points (car, truck, other-vehicle) into $S_{assoc}^{b}$, whereas classes with naturally small point counts (person, bicyclist, motorcyclist, motorcycle, bicycle) are grouped into $S_{assoc}^{s}$. As shown in Figure 6, ICP-4D excels at maintaining instance consistency not only for vehicle classes but also for challenging small object classes, outperforming the baselines. Additionally, it is worth noting that all classes except bicycle exhibit favorable association performance (see Figure 7). This is attributed to the robustness of ICP-4D under scenarios where object sizes and occurrence frequencies vary significantly.
Figure 6: Comparison of association quality by size group.
Figure 7: Class-wise association performance.

### v2

@@ 3.2 Geo-4D
$j^{*}(i)={\underset{j}{\operatorname*{arg\,max}}}\hat{\mathcal{Q}}_{ij},\quad\mathcal{C}^{(k)}_{\text{GGSM}}=\{(p_{i},q_{j^{*}(i)})\mid i=0,\dots,I-1\}.$ (9)
That is, we approximate point-to-point correspondences by selecting the highest transport probability from the minimized transport plan. As a result, this strategy is based on considering the instance instead of local points (see Figure 2), thereby alleviating the vulnerability to outliers and incorrect predictions.
Returning to the ICP procedure, we follow the standard process, initializing the transformation with a histogram-based method [lin2024icp] to reduce sensitivity to initial estimates and subsequently running iterative refinement to get the final rigid transformation. After obtaining the transformation ($\mathbf{R},\mathbf{t}$), we determine whether two instances correspond by evaluating the IoU between their transformed point sets. Specifically, points with errors smaller than the predefined threshold $\tau_{dist}$ are treated as inliers. Then, in order to filter valid matching candidates, we evaluate alignment accuracy between source and destination instances using IoU scores as follows:
$\delta(a^{\prime\prime}_{\text{src}},b^{\prime\prime}_{\text{dst}})=\mathds{1}[\text{IoU}(\mathcal{F}(\mathcal{P}(a^{\prime\prime}_{\text{src}})),\mathcal{P}(b^{\prime\prime}_{\text{dst}}))\geq\tau_{\text{iou}}],$ (10)
where $\text{IoU}(\cdot)$ denotes the IoU calculation, and $\mathcal{F}$ applies the estimated transformation obtained by ICP. Using Eq. 10, we assign the same instance ID to matched candidates. In our experiments, this simple matching strategy consistently outperforms more complex assignment schemes. Additional comparisons are provided in the supplementary material.
Missing Instance. Occlusions stemming from the inherent limitations of LiDAR sensor often break temporal continuity beyond two consecutive scans. To mitigate this, we introduce a memory bank $\mathcal{B}$ that caches the world coordinates, semantic labels, and instance IDs of unmatched objects in both static and dynamic association stages. Owing to bounded memory capacity, $\mathcal{B}$ keeps only entries observed within a short window of the previous $w_{\text{mem}}$ scans. If $\mathcal{B}$ is not empty, we retrieve the stored instance information and apply the global geometric association strategy to restore temporal consistency.
@@ 4 Experiments
4.1 Experiment Setup
Datasets. We choose two representative benchmarks. SemanticKITTI [behley2019semantickitti] captures point clouds using a 64-beam LiDAR at 10 Hz and provides point-wise annotations for 19 semantic classes (8 Things / 11 Stuff). In contrast, panoptic nuScenes [fong2022panoptic] uses a 32-beam LiDAR operating at 2 Hz and offers point-wise annotations for 16 semantic classes (10 Things / 6 Stuff). This contrast serves as a valuable testbed for assessing generalization across dense (SemanticKITTI) and sparse (nuScenes) LiDAR settings.
Evaluation Metrics. To evaluate both tracking and segmentation quality, we adopt the widely used LSTQ metric [Aygun_2021_CVPR], defined as $\text{LSTQ}=\sqrt{S_{cls}\times S_{assoc}}$. Here, $S_{cls}$ measures how accurately semantic labels are assigned and $S_{assoc}$ evaluates how consistently instance identities are maintained throughout the sequence, independently of semantic accuracy. However, $S_{assoc}$ is computed only from true positive associations weighted by IoU scores. Thus, when prediction masks are stable and yield high IoU scores, ID switch causes a sharper drop in this score. Consequently, while Geo-4D primarily influences $S_{assoc}$, the LSTQ must be taken into account as a meaningful indicator of temporal consistency.
Implementation Details. Our proposed Geo-4D can be incorporated into any 3D LiDAR panoptic segmentation network. For fair comparison, we basically follow the configurations of two baselines, DS-Net [hong2021lidar] and Mask4Former [yilmaz2024mask4former]. Since Mask4Former is originally developed for 4D panoptic segmentation, we train it using the official settings with the input modified to a single scan instead of multiple sweeps. For instance association, thresholds for filtering instance pairs are set to $\tau_{center}=\tau_{cov}=\tau_{dist}=0.1$ and $\tau_{iou}=0.2$. Additionally, the short window $w_{\text{mem}}$ is set to $3$ for missing instances, the regularization parameter for optimal transport is set to $\epsilon=0.2$, and the number of ICP iterations is fixed to 30. All experiments are conducted on a single RTX A6000.
4.2 4D LiDAR Panoptic Segmentation Results
SemanticKITTI. Table 4.2 reports the experimental results on SemanticKITTI validation and hidden test set. Here, we clearly observe two insights. First, Geo-4D achieves the highest LSTQ score across all data splits regardless of the number of input scans. Specifically, while (a) IoU-based association approaches rely on 22-44 more input point clouds, our Geo-4D achieves superior performance using only a single scan, showing improvements of up to 13.2 percent points in $S_{assoc}$ and 10.0 percent points in LSTQ. Furthermore, by comparing with the learning-based approaches (b) and (c), another noticeable aspect is that Geo-4D consistently outperforms recent state-of-the-art methods on both the validation and test sets, whereas Mask4D and CA-Net drop in performance on the test set. This highlights the robustness of our direct geometric registration with raw point clouds, rather than relying on training-dependent feature association.
Table 1: Quantitative results on SemanticKITTI [behley2019semantickitti]. We highlight the best scores in bold and the underline indicates the second-best scores. Note that * represents that the results are obtained from our implementation following the official training setup.
@@ 4.1 Experiment Setup
4D-PLS [Aygun_2021_CVPR] 4 62.7 65.1 60.5 65.4 61.3 56.9 56.4 57.4 66.9 51.6
4D-StOP [kreuzberg20224d] 4 67.0 74.4 60.3 65.3 60.9 63.9 69.5 58.8 67.7 53.8
Eq-4D-StOP [zhu20234d] 4 70.1 77.6 63.4 66.4 67.1 67.0 72.0 62.4 69.1 60.9
Mask4Former [yilmaz2024mask4former] 2 70.5 74.3 66.9 67.1 66.6 68.4 67.3 69.6 72.7 65.3
@@
KPConv [thomas2019kpconv]+PP [lang2019pointpillars]+MOT [weng20203d] 1 46.3 37.6 57.0 64.2 54.1 38.0 25.9 55.9 66.9 47.7
KPConv+PP+SFP [mittal2020just] 1 46.0 37.1 57.0 64.2 54.1 38.5 26.6 55.9 66.9 47.7
black!10black DS-Net    [hong2021lidar] 1 - - - 63.5 64.6 62.0 - - 60.6 66.9 52.0
@@
Table 2: Quantitative results on panoptic nuScenes [fong2022panoptic].
@@
4D-StOP [kreuzberg20224d] 4 60.5 62.5 58.6 74.4 54.9
Eq-4D-StOP [zhu20234d] 4 67.3 73.7 61.5 76.4 58.7
Mask4Former [yilmaz2024mask4former] 2 67.6 63.0 72.5 76.6 66.3
(b) Query-propagated Association
Mask4D [marcuzzi2023mask4d] 1 58.5 54.1 63.1 75.5 55.7
(c) Detect & Track Association
@@
+ CA-Net 1 63.4 56.6 71.0 76.0 65.9
+ Geo-4D (Ours) 1 67.9 65.0 71.0 76.0 65.9
Table 3: Comparisons across multiple metrics on SemanticKITTI validation set. ( ) denotes the number of input scans.
Method LSTQ MOTSA sMOTSA PTQ sPTQ
4D-StOP (2f) 66.4 31.0 27.1 53.4 53.7
EQ-4D-StOP (2f) 68.9 43.6 39.6 56.4 56.7
Mask4D (1f) 71.4 52.1 48.1 58.4 58.5
Mask4Former (2f) 70.5 47.3 42.1 48.3 48.5
Mask4Former (1f) - - - - -
+ CA-Net 70.4 48.4 43.1 48.9 49.1
+ Geo-4D (Ours) 72.7 73.3 65.7 59.6 59.7
@@
nuScenes. Table 3 shows the 4D panoptic results on panoptic nuScenes dataset. Consistent with the observations on the SemanticKITTI, Geo-4D delivers notably stronger instance association performance, outperforming existing methods in terms of LSTQ. It is important to note that baselines generally suffer when the LiDAR scans are captured with a long interval. In such scenarios, multi-scan aggregation benefits from denser temporal support and thus exhibit strong performance. For example, EQ-4D-StOP with 4 aggregated scans surpasses Mask4Former with 2 scans by a substantial margin in the association metric. Nevertheless, Geo-4D demonstrates that even without accessing multiple scans, a single-scan input is sufficient to achieve competitive performance. This result implies that Geo-4D remains highly effective even under sparse LiDAR configurations.
Results on Multiple Metrics. We further present additional results with multiple metrics to enable a more comprehensive comparison. MOTSA [voigtlaender2019mots] is widely used for evaluating multi-object tracking and segmentation. Although MOTSA emphasizes the recognition quality rather than temporal consistency, it nonetheless offers a straightforward indicator of the ID switches. Moreover, PTQ [hurtado2020mopt] is designed to jointly evaluate panoptic segmentation and tracking quality in a holistic manner. In Table 3, evaluations are carried out on the SemanticKITTI [behley2019semantickitti] validation set. We observe that Geo-4D consistently surpasses all baselines across multiple metrics, including LSTQ. This indicates that the strength of Geo-4D goes beyond any particular metric, offering a robust and practically reliable solution for instance association.
Figure 3: Qualitative comparison on SemanticKITTI validation set. We visualize the association results over five consecutive scans for both the baselines and our Geo-4D. Different colors represent different instances. The dotted boxes zoom in for clear comparison. For clarity, we use circled numbers (①–⑤) to represent the frame indices.
@@
Comparison of Efficiency. Next, we examine computational efficiency in terms of memory usage and runtime. For each method, we report the peak GPU memory consumption and the average runtime measured over the entire validation set. Figure 4 illustrates the results, supporting that Geo-4D offers the most well-balanced trade-off between performance and computational overhead. First, when compared with the IoU-based association approaches, we observe that Geo-4D requires 60.7%–79.4% less memory and achieves a 26.1%–45.6% runtime efficiency improvement, all while delivering superior LSTQ performance. While Mask4D exhibits high efficiency in runtime, it inevitably incurs substantial memory overhead due to its transformer-based architecture design. Lastly, CA-Net is computationally efficient; however, it fails to obtain reliable association performance. This gap is already observable on the validation set and becomes even more pronounced on the hidden test set, as highlighted in blue in Table 4.2.
Figure 4: We illustrate the efficiency trade-offs between memory usage (left) & runtime (right) and performance across different methods.
@@
Qualitative Analysis. We deeply analyze the effectiveness of Geo-4D qualitatively by presenting the visualization results in Figure 3. In the first row, there are two bicyclists riding in a line. Both ground truth and Geo-4D consistently distinguish the two bicyclists across time. However, Mask4Former struggles to preserve identity consistency when objects are spatially adjacent or overlapping due to its IoU-based association. Furthermore, Mask4D employs learnable tracking queries for association, which can lead to confused tracking queries when objects share similar visual or spatial features. Then, in the second row, we emphasize the effectiveness of our memory bank in handling missing instances. Even when the 3D panoptic model fails to detect the person colored in green at the second time step, the memory bank successfully retrieves the missing instance and preserves its identity across scans. In contrast, Mask4Former experiences ID switching when it fails to detect an instance due to the absence of overlapping regions. Consequently, these observations underscore the importance of explicit geometric matching in our approach and a well-structured association pipeline.
@@
4.3 Ablation Study
We conduct a series of ablation studies to validate the effectiveness of each component in our framework. All experiments are performed on the SemanticKITTI [behley2019semantickitti] validation set, with Geo-4D integrated into Mask4Former [yilmaz2024mask4former].
Module Design. Table 4 shows the effects of our proposed methods. Exp. #1 directly applies the vanilla Iterative Closest Point (ICP) for associating instances, serving as a baseline to examine the feasibility of ICP in an unsupervised manner. Despite its strong matching capability, ICP is highly sensitive to the quality of initial alignment. In Exp. #2, we observe a little performance gain when applying histogram-based initialization proposed by [lin2024icp], which stabilizes the initial alignment. Notably, our Global Geometry-aware Soft Matching (GGSM) brings further improvements without the influence of initial pose (see Exp. #3). The key takeaway is that instance-aware correspondence to alleviate the instability of local point matching is highly beneficial for obtaining reliable matching points between noisy instance segments, and as a result helps to estimate an accurate transformation. Exp. #4 demonstrates that coupling these components yields a complementary effect, leading to more robust instance association and improved overall performance. Lastly, Exp. #5 demonstrates the effectiveness of incorporating a memory bank for handling occluded and missing instances, resulting in superior performance.
@@ 4.3 Ablation Study
Effect of Different Types of Instances. To validate our carefully designed pipeline with respect to different instance types, Table 5 presents the association results and average inference time per a single scan. First, we employ a statistics-based geometric filtering for static instances to match candidates over all detected objects. Despite requiring minimal computation, only a small subset of instances remain consistently tracked, 18.5 percent points indicating that static cues alone are insufficient for reliable association. In contrast, relying solely on global geometry association achieves high performance, but is computationally expensive. Thus, combining these strategies reduces the runtime by 48.2% while simultaneously improving performance, as it significantly narrows the matching candidates for global geometry association. Consequently, our full Geo-4D enhances the tracking continuity without sacrificing computational efficiency.
Hyperparameter Analysis In this section, we examine the impact of different matching thresholds on instance association performance. We empirically determine the matching thresholds $\tau_{iou}$, $\tau_{dist}$, $\tau_{center}$, $\tau_{cov}$} over {0.1, 0.2, 0.3, 0.4, 0.5}. Here, we observe three insights. First, extremely low value of threshold $\tau_{iou}$ often yields decent performance due to over-association, whereas high thresholds lead to over-filtering and degraded matching quality. Second, excessively large $\tau_{center}$ and $\tau_{cov}$ cause static instances to be matched too liberally, leading to unreliable association. Lastly, our proposed Geo-4D delivers performance gains over all baselines, regardless of the hyperparameter choices.
Figure 5: We illustrate the performance variation w.r.t. the hyperparameters of Geo-4D. The dashed line marks Mask4D, the strongest-performing baseline.
Table 6: DBSCAN instance refinement.
@@
Analysis of DBSCAN Instance Refinement. In this specific experiment of integrating Geo-4D with Mask4Former, we apply DBSCAN clustering [campello2013density] as done in the original Mask4Former setting. Table 6 summarizes the results, where we derive two major insights. First, Geo-4D consistently surpasses existing state-of-the-art methods, both with and without the additional clustering step, achieving at least a 0.4 percentage point gain in LSTQ. Furthermore, we discover that the increase of runtime primarily results from the post-processing step. Without this step, Geo-4D delivers the most efficient runtime of 840 ms compared to all other methods, with only a marginal loss in performance. These observations clearly highlight that Geo-4D offers both high efficiency and strong performance, and can further benefit from improvements in the backbone network.
@@
Effect of Global Geometry-aware Soft Matching. The Global Geometry-aware Soft Matching (GGSM) forms instance-aware point correspondence between source and target points. Owing to this strategy, accurate error measurements can be obtained even in the presence of outliers, as the estimation is not dominated by local proximity. Figure 6 visualizes the point-wise correspondences obtained from two consecutive scans, with the current scan serving as the source and the previous scan as the target. For this analysis, we extract each point segment at the initial stage of ICP, as this stage retains its geometric structure intact. As shown in the highlighted region, the nearest-neighbor matching is biased toward the closest point, ignoring the underlying point-wise relationships of the instance. In contrast, applying GGSM on the baseline maintains the geometric structure, yielding correspondences that align consistently with the instance’s motion direction.
Figure 7: (a) Comparison with representative baselines across different association strategy categories on rigid and non-rigid object classes. (b) Class-wise performance with respect to object class types and the average number of points per instance.
Rigid vs. Non-Rigid Class Analysis. We categorize object classes into rigid (car, truck, other-vehicle, bicycle, motorcycle) and non-rigid (bicyclist, motorcyclist, person) groups based on their physical characteristics. As shown in Figure 7(a), Geo-4D outperforms representative IoU-based and query-based association baselines across both rigid and non-rigid classes. Furthermore, Geo-4D consistently maintains strong association performance across all object classes (see Figure 7(b)). These results are attributed to the robustness of our proposed global geometry-aware association strategy, which captures holistic geometric structure and enables reliable association under partial non-rigid motion and significant variations in object size.